(1 point) Find d2ydx2 for
y=csc(x)
d2ydx2=
1
Expert's answer
2020-02-26T10:41:02-0500

"y=csc(x)=1\/sin(x)=sin^{-1}(x)"


using "(f(g(x)))'=f'(g(x))*g'(x)"


"y'=(-1)*sin^{-2}(x) *cos(x)"


using "f(g(x))'=f'(g(x))*g'(x)" and "(fg)'=f'g+g'f"


"y''=(-1)* [ (-2)sin^{-3}(x) *cos(x)*cos(x) + sin^{-2}(x) *(-sin(x))]"


simplify:


"y''=2cot^2(x)csc(x)+csc(x)"


alternate form:


"y''=csc(x) (cot^2(x) + csc^2(x))"